OpenStudy (anonymous):

SOMEONE PLEASE HELPP ME!!!!! A regular pentagon has sides measuring 8 cm. The pentagon is enlarged using a scale factor of 5/2 Part 1) What are the lengths of each side of the enlarged pentagon? Part b) What is the perimeter of the enlarged pentagon?

jimthompson5910 (jim_thompson5910):

"The pentagon is enlarged using a scale factor of 5/2 " translation: each side or dimension is multiplied by 5/2 5/2 = 2.5

jimthompson5910 (jim_thompson5910):

so let's say each side is 6 cm long you multiply each side by 5/2 or 2.5 to get 6*(5/2) = 6*(2.5) = 15 each side is now 15 cm long. This is not the answer as it's just an example

OpenStudy (anonymous):

Ok thank you @jim_thompson5910 this should help me :)

jimthompson5910 (jim_thompson5910):

I'm glad it will.

OpenStudy (anonymous):

Hey @jim_thompson5910 can you help me find the perimeter please

jimthompson5910 (jim_thompson5910):

how long is each side after you do the scaling?

OpenStudy (anonymous):

I got 20cm

OpenStudy (ilovecake):

If you are done with this. then close this question

jimthompson5910 (jim_thompson5910):

each side is 20 cm there are 5 sides (this is a pentagon) each side is congruent or equal to one another (it's a regular polygon)

jimthompson5910 (jim_thompson5910):

you just add up all the sides to get the perimeter (each side is 20 cm)

OpenStudy (anonymous):

Oh ok thank you so much @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem